OpenStudy (anonymous):

solve the following system of equation by graphing and finding the intersection point 2X-2Y=-8 & 4X+2Y=-4 PLEASE BREAK IT DOWN

OpenStudy (anonymous):

Give me a sec and I can help.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2x-2y=-8 Here, we want Y by itself. So first we must subtract 2x on each side. This yields -2y=-8-2x Now, to get Y by itself, we have to divide each side by -2. This yields: y=4+x or rearrange to get: y=x+4 which is in y=mx+b form and you can graph. You should be able to do the second equation on your own the same way. Graph both and find the point at which they intersect/cross each other/are equal.

OpenStudy (anonymous):

2,6

OpenStudy (anonymous):

just graph them one by one

OpenStudy (anonymous):

so then u dont do the y multiplying both equations]

OpenStudy (anonymous):

like y= 2 and x=1 plus 1 y=x+1

OpenStudy (anonymous):

for the second one i got y=-2-x

OpenStudy (anonymous):

I put both of them in slope-intercept form, then graphed them, I got (2,6) in both so the intersection is (2,6)

OpenStudy (anonymous):

i have a problem idr wat to do after i put it in y=mx=b

OpenStudy (anonymous):

i got y=2x+2

OpenStudy (anonymous):

oooo i wrote the wrong number

OpenStudy (anonymous):

just make sure their both equal on the side like y=2x+2 y can be 6 and the x in the 2 can be 2 so 2x2 =4+2=6 6=6

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thank yiu

OpenStudy (anonymous):

x=-2 , y=2